APPENDIX A

Changes for version 2.01

• Printing raster pictures (IFF ILBM files and PostScript) in landscape orientation:

If you own a PCL-5 compatibile printer (e.g. HP LaserJet 4), you don’t have to manually rotate picture to be printed correctly in landscape orientation. If your printer doesn’t support PCL-5, you will still have to rotate picture manually.

• Added two new options for command line parameters/tooltypes:

X=XOFFSET and Y=YOFFSET let you specify the position of the top left corner. Default is 0, which means that this corner is 1 inch from the left edge and 1 inch from the top edge of the paper. If value for X/Y offset is positive, the corner will move right (down), if it is negative, it will move left (up).

X and Y offset should be specified in 1/300” (300 = 1”, 150 = 1/2”, etc.).

• Resolution problems fixed:

Printing at resolution other than 300 dpi gave corrupted output.

• When the user cancelled the requester, the program didn’t always free its resources — bug fixed.

Special thanks to Matej Skobe and Thomas Ganter for bug-reports and suggestions.